write all possible competitions of length and breadth for a rectangle with area 36 sq cm. also find perimeter in each case
Answer by Alan3354(69443)

write all possible competitions of length and breadth for a rectangle with area 36 sq cm. also find perimeter in each case
--------------
Maybe you mean combinations.
-------
Area = L*W = 36
L = 36/W
--------
Perimeter = 2L + 2W
P = (72/W) + 2W